g(t)=3t/e^6t

f(x)= e^x sine^x



1
Expert's answer
2021-01-11T18:56:52-0500

Here we have "g(t)=\frac{3t}{e^{6t}} \\"


So, differentiating both sides:-



"g(t)=\frac{3t}{e^{6t}}\\~\\
\Rightarrow g'(t)=\frac{e^{6t}(3)-3t(6e^{6t})}{e^{12t}}\\~\\
\Rightarrow g'(t)=\frac{3e^{6t}-18te^{6t}}{e^{12t}}\\~\\
\Rightarrow g'(t)=\frac{3-18t}{e^{6t}}"


And also we have "f(x)=e^xsin(e^x)"


So, differentiating both sides:-



"f(x)=e^xsin(e^x)\\
\Rightarrow f'(x)=sin(e^x)e^x+e^xcos(e^x)e^x\\
\Rightarrow f'(x)=e^xsin(e^x)+e^{2x}cos(e^x)"


So, we have our derivatives of the two given functions.